I am having trouble with finding the six trigonometric functions of radians which are not commonly found on the unit circle( for example 17pi/2, 27pi/4, 15pi/6). I understand how to find these six functions with certain radians or degree measures common to the unit circle, but no others. Thank you for your time.
Answer by lwsshak3(11628)

I am having trouble with finding the six trigonometric functions of radians which are not commonly found on the unit circle( for example 17pi/2, 27pi/4, 15pi/6).
..
17π/2=8π+π/2= rotating counter-clockwise eight π radians or four 2π radians plus another π/2 radian.
So, the trig functions are evaluated for(π/2) radians.
sin(π/2)=1
cos(π/2)=0
tan(π/2)=ud
csc(π/2)=1
sec(π/2)=0
cot(π/2)=0
..
27pi/4= =6π+3π/4= rotating counter-clockwise six π radians or three 2π radians plus another 3π/4 radian.
So, the trig functions are evaluated for 3π/4) radians.
...
15pi/6=2π+π/2= rotating counter-clockwise two π radians or one 2π radians plus another π/2 radian.
So, the trig functions are evaluated for π/2 radians.